If (x – 2) is a factor of P(x)= ax³ +3x² - 2x + a , determine the value of a
1
Expert's answer
2012-04-19T10:54:41-0400
if( x-2) is a factor of P(x), then P(2)=0.
P(2) = 8*a + 3*4 - 2*2 + a =
0.
Then solve the equation 8a +12 - 4 + a = 0
9a + 8 = 0
a = -8/9